What mass of nitrogen is needed to react with 21.25 g of hydrogen?


1
Expert's answer
2022-05-11T01:33:05-0400

N2 + 3H2 = 2NH3

m(N2) = (m(H2)•Mr(N2)) / 3Mr(H2)

m(N2) = (21.25 g • 28 g/mol) / 3 • 2 g/mol = 99.17 g